# ORCA: An Open-Source, Reliable, Cost-Effective, Anthropomorphic Robotic Hand for Uninterrupted Dexterous Task Learning

### v1

@@ III-E Tactile Sensing
However, wear and tear on the silicone skin can drastically affect the AT of the sensors. Although the sensor mounted on the ring finger of the hand showed no degradation after thousands of grasp cycles as part of the experiments described in Section III-A, the AT for the sensor mounted on the pinkie finger increased to $6.38N$ due to the degradation of the silicone skin creating an air gap between the silicone skin and the mounted FSR sensor.
IV Additional Results and Discussion
@@
IV-A Teleoperation - Dexterous Manipulation
  • Grab a plush toy (about the size of 3 large cubes) from the table top.
  • Grab a tennis ball lying on the table
  • Spin a fidget toy for $2\,s$. The fidget spinner is placed on a finger by hand, but the grasping and spinning are done purely by teleoperation.
  • Pick up a piece of paper lying on a shut box.
  • Pick up a cup and pour its contents ($50\,ml$ water) into another cup.
IV-B Imitation Learning - Repeated Pick & Place
To highlight the reliability of both the hardware platform and the trained IL policy, we designed a continuous pick-and-place evaluation task. The robotic hand is required to pick up a cube (6 cm in side length) from a table and place it on a sliding surface, which then causes the cube to fall back onto a random location on the table (Figure 5, Figure 6-A).
To evaluate different policies, we collected the ratio of failure positions to success positions (regarding picking up the cube) within a testing area for 60 iterations (10 per sub-area), as shown in Figure 7. The most successful policy, using only masked images of the cube, is deployed for 7h 17min (approx. 2,000 grasping cycles) with no human intervention on the ORCA hand’s hardware and minimal intervention in aiding in the pick-and-place task (Figure 1-B). Throughout the test, the policy maintained consistent performance, with no tendon slack or rupture, and the experiment was concluded not due to failure but because it sufficiently demonstrated the system’s reliability and effectiveness. For extended videos and time-lapses of the reliability test, please refer to the ORCA project website.
IV-C Tactile Sensing
IV-D Ease of Assembly
The ORCA hand is designed to be easily assembled and repaired. Given all the necessary mechanical parts, it takes one person less than 8 hours to assemble one entire ORCA hand with integrated sensors, tensioned tendons, and soft skin. A detailed step-by-step instruction on how to assemble and repair the hand will be available on the project website.
@@ IV-D Ease of Assembly
Total $50$% $86.\overset{\_}{6}$$70$%
Figure 7: Testing area with respective policy success rates (out of 10)
@@ V Conclusion
The ORCA hand offers an accessible solution for advancing robotic manipulation, and we hope to have laid the foundation for real-world tasks and cutting-edge manipulation research. A great deal of work has gone into making this human-like, compliant hand as robust and reliable as possible while remaining versatile, easy to control, and cost-effective.
Nevertheless, our study has limitations that suggest directions for future research. Prolonged use of the ORCA hand necessitates manual re-tensioning to sustain optimal performance. To address this, we aim to develop a re-tensioning mechanism that autonomously reduces tendon slack without the need of human intervention. This advancement could enable even longer operation times, potentially accelerating real-world RL applications for dexterous manipulation research. Future work on our hand may also include full integration of the sensors into the learning pipeline, as well as more advanced deep learning algorithms to learn complex interactions with the environment more effectively.
@@ ACKNOWLEDGMENT
Thanks to Robert Jomar Malate and the Real World Robotics course team at ETH Zurich who initialized and greatly enabled this project. We also thank Carlo Sferrazza and the Berkeley Robot Learning Lab for providing access to their LEAP Hand for comparison.
The authors are grateful for funding through a research collaboration with armasuisse. Filippos Katsimalis thanks the Bodossaki Foundation, SYN-ENOSIS, and the Onassis Foundation for their invaluable financial assistance. Aristotelis Sympetheros thanks the John S. Latsis Public Benefit Foundation and SYN-ENOSIS for supporting his postgraduate studies.

### v2

@@ III-E Tactile Sensing
Total 5050% $86.\overset{\_}{6}$% 7070%
Figure 7: Testing area with respective policy success rates (% out of 10)
However, wear and tear on the silicone skin can drastically affect the AT of the sensors. Although the sensor mounted on the ring finger of the hand showed no degradation after thousands of grasp cycles as part of the experiments described in Section˜III-A, the AT for the sensor mounted on the pinkie finger increased to $6.38N$ due to the degradation of the silicone skin creating an air gap between the silicone skin and the mounted FSR sensor.
IV Additional Results and Discussion
IV-A Teleoperation - Dexterous Manipulation
  • Grab a plush toy (about the size of 3 large cubes) from the table top.
  • Grab a tennis ball lying on the table
  • Spin a fidget toy for 2s2\,s. The fidget spinner is placed on a finger by hand, but the grasping and spinning are done purely by teleoperation.
  • Pick up a piece of paper lying on a shut box.
  • Pick up a cup and pour its contents (50ml50\,ml water) into another cup.
IV-B Imitation Learning - Repeated Pick & Place
To highlight the reliability of both the hardware platform and the trained IL policy, we designed a continuous pick-and-place evaluation task. The robotic hand is required to pick up a cube (6 cm in side length) from a table and place it on a sliding surface, which then causes the cube to fall back onto a random location on the table (Figure˜5, Figure˜6-A).
To evaluate different policies, we collected the ratio of failure positions to success positions (regarding picking up the cube) within a testing area for 60 iterations (10 per sub-area), as shown in Figure˜7. The most successful policy, using only masked images of the cube, is deployed for 7h 17min (approx. 2,000 grasping cycles) with no human intervention on the ORCA hand’s hardware and minimal intervention in aiding in the pick-and-place task (Figure˜1-B). Throughout the test, the policy maintained consistent performance, with no tendon slack or rupture, and the experiment was concluded not due to failure but because it sufficiently demonstrated the system’s reliability and effectiveness. For extended videos and time-lapses of the reliability test, please refer to the ORCA project website.
IV-C Tactile Sensing
@@ V Conclusion
The ORCA hand provides an accessible platform for advancing robotic manipulation, aiming to support real-world tasks and cutting-edge research. It is designed to be human-like, compliant, robust, versatile, easy to control, and cost-effective.
Despite these strengths, limitations remain. Prolonged use requires manual re-tensioning to maintain performance. To overcome this, we plan to develop an autonomous re-tensioning mechanism to reduce tendon slack without human intervention, enabling longer operation and supporting real-world RL applications. Future work also includes integrating sensors directly into the learning pipeline and applying more advanced deep learning methods to handle complex environmental interactions.
@@ ACKNOWLEDGMENT
We thank Robert Jomar Malate and the Real World Robotics course team at ETH Zurich for initiating and enabling this project, and Carlo Sferrazza and the Berkeley Robot Learning Lab for access to their LEAP Hand. This work was funded through a research collaboration with armasuisse. Filippos Katsimalis acknowledges support from the Bodossaki Foundation, SYN-ENOSIS, and the Onassis Foundation. Aristotelis Sympetheros acknowledges support from the John S. Latsis Public Benefit Foundation and SYN-ENOSIS.